OpenStudy (anonymous):

. Which fraction is not equivalent to 30/48 is it 5/8, 15/24, 20/38 or 35/56

OpenStudy (kaylieanime):

Well one way to help is to reduce everything. That makes it easier to compare.

OpenStudy (anonymous):

30/48 divide above an below by 2 = 15/24, so they are equivalent 35/56 divide above an below by 7 = 5/8, so they are equivalent 15/24 divide above an below by 3 = 5/8, so they are equivalent but 20/38 is not equivalent